OpenStudy (anonymous):

Stan is making a trail mix that has a combination of peanuts and raisins. The mix is 3 parts peanuts and 1 part raisins. How many cups of raisins does he need to use for a 16-cup recipe?

OpenStudy (anonymous):

mixture is=peanuts:raisins=3:1 so in (3+1)cup i.e 4 cup there will be 1 cup raisin so for 16 cup recipe raisin required is=(1/4)*16=16/4=4 cup. So your answer is 4 cups